OpenStudy (ybballerrah908):

An electrician charges a flat fee plus an hourly rate for a service call. She charges a total of $94 for a 3-hour service call on Monday and $130 for a 5-hour service call on Tuesday. How much should she charge for a 2-hour service call?

OpenStudy (robtobey2):

Let f and t be the flat fee and hourly rate respectively. Solve the following for f and t: {f + 3 t = 94, f + 5 t = 130}

OpenStudy (mathmate):

Electricians usually charge fixed rate travelling + an hourly rate. If he charged $94 for a 3 hour call, and $130 for a 5 hour call, then the charge for the two extra hours is $(130-94)=$36, or his hourly rate is $18. For a two hour call, I would just take off $18 from the 3 hour call, or $(94-18)=$76. Alternatively, we can work out the fixed travelling charge, which is $94 less 3 hours, or $(94-3*18)=$(94-54)=$40. Then for any number of hours x of the call, we have the formula $(40+18x) for his charge. But this is not what the question wants.